OpenStudy (anonymous):

3square root 6 multiplied by 6 is???

OpenStudy (johnweldon1993):

You would just do 3 * 6....and bring the √6 along

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

idk how to get teh square root sign when i type the question in

OpenStudy (johnweldon1993):

well I used an alt key....you can look it up and figure out how to use them do you understand the answer though?

OpenStudy (anonymous):

no well um i got 6 square root 3 for the answer idk if its right?

OpenStudy (johnweldon1993):

not quite \[3\sqrt{6} * 6\] \[3 * 6 = 18\] answer is \[18\sqrt{6}\]

OpenStudy (anonymous):

\[6\sqrt{3 }\]

OpenStudy (anonymous):

how u get 18 sq root 6?

OpenStudy (johnweldon1993):

like I said.....when you have a number...with a square root....being multiplied by another number....you multiply the 2 numbers....and just ignore the square root until the end so with your 3√6 times 6 you multiply the 3 and the 6 to get 18 and you bring back the √6 for the answer 18√6

OpenStudy (anonymous):

\[3\sqrt{6\times \sqrt{6}}\] is my question

OpenStudy (anonymous):

well then thanks anyways

OpenStudy (anonymous):

i thought u have to look for factors of 6

OpenStudy (anonymous):

i mean 36